10–53. Determine the moment of inertia of the area about axis. the y у 3 in.3 in.-- 6 in. 2 in. 4 in. Probs. 10-52/53
Answer #1

So lets label our figure and set it up as follows

10–53. Determine the moment of inertia of the area about the y axis. у 3 in.3 in.-- D B 6 in. 0 E 2 in. 4 in. X А F Probs. 10

We will calculate moment of intertia as follows

Momement of inertia of blue shaded area = Moment of intertia of big rectangle ABCG + Moment of inertia of Triangle COE + Moment of Inertia of rectangle GOEF- Moment of inertia of circle

So we have the areas as follows. Use -ve notation circle

Part Area Type Area formula Area (in) 1 2 3 3 Rectangle ABCG Triangle COE Rectangle GOEF Circle AB * BC 1/2 *CO * DE GO *OE p

We can then calculate the centroids as follows

Part Area Type Area formula Area (in?) Centroid x cooordiate Centroid y cooordiate 1 1.50 2 3 3 Rectangle ABCG Triangle COE R

For rectangle ABCG Iyc = 1/12 * 10 * 63 = 180

For triangle COE Iyc (of Triangle COE) = bh3/36 = 1/36* 3 * 63 = 18

For rectangle GOEF Iyc (of rectangle GOEF) = 1/2 * 4 * 33 = 54

For circle Iyc of circle = pi* radius4/4 = 22/7 * 24/4 = 12.571

So we have the following

Part Area Type Area formula Area (in) Centroid x cooordiate Centroid y Area * Area * cooordiate Centroid(x) Centroidly) Momen

by parallel axes theorem we know that

Iy = Iyc + (dx)2 *Area

so we have

Part Area Type Area formula Area (in) Centroid x cooordiate Centroid y Area * Area * cooordiate Centroid(x) Centroidly) Momen

So moment of inertia of the shaded area = 800.428

=====================================================================================================

Please rate thumbs up to suppor and encourage us to keep giving good solutions.